Meeting Information

  • 1600 UT (1000 MDT)
  • AOC 317; IP: 146.88.16.8; phone: +1 575-835-7370

Agenda/Minutes

News, Travel, Visitors (McMullin)

  • Overall status: In DnC configuration; begin move to C config on 04Oct. Preliminary testing of time averaging looks good for C config deployment (see later talk by Emmanuel). Good progress on the 4-band testing (properly identified the band of operation, after identifying the misbehaving antennas (with only 3.5 MHz BW). 3-bit sampler testing narrowing. Correlator BlB HW failure; replaced. Several RSRO test observations over the last week. Spending today on science.
  • Key dates:
    • 13-22 Sep: 74 MHz array testing campaign
    • 15 Sep: Call for Proposals
      • Note: Ku/X-wide will not be offered in main call (allowed for exploratory).
    • 01 Oct: Proposal deadline (A config; lowband for end of B config)
    • 04 Oct: Begin move to C configuration
    • 13-15 Oct: Phased Array testing campaign
  • Staff Travel:
    • Rupen: 11 Sep - 21 Sep
    • Sowinski: 20 Sep - 15 Oct
    • McMullin: 01-10 Oct (vacation/ALMA)
    • van Moorsel: 11-15 Oct
    • Wrobel: 24 Oct-27 Oct
    • DRAO: Dougherty: 11 Oct - 12 Oct (SKA; ABQ)
  • Visitors:
    • Brent Carlson: 18 October - 25 October; second trip post SKA meeting TBD
    • DRAO: Dougherty: 12 Oct - 17 Oct (NRAO-DSOC)
    • Adam Deller: 09 Oct - 16 Oct (phased array testing)
    • Joe Lazio: 12-15Oct (phased array testing)

Hardware status and updates (Langley, Perley, Rupen)

  • Array (Perley/Langley):
    • Notes: 4-band dipoles coming off on Wednesday. EA14 multiple receivers (out of the barn); EA06 (L band); EA24 (Ku band).
      • Antennas with 3-bit samplers: EA12, EA15, EA22, EA28
      • Antennas with 4-band dipoles: EA01(West), EA03(East), EA06(North), EA11(East), EA19(West), EA28(North)
    • Links:
  • WIDAR (Rupen):
    • Notes: HW failure on Friday with one BlB; replaced yesterday (not fully connected-follow-up today); tested quadrant 2 BlBs and using this partially now (connectivity more fully established with BC's visit in October).
  • Computing (Robnett):
    • Notes: Lustre file system disks will be made available; policies still being discussed but this will bring 60 TB onto the table. Using a subset as testing group. New version of MCAF needs testing. Following up on missing/mislabeled scans (varies substantially from project to project; missing data seems to be a small effect but mislabeled scans can occur frequently in small files. Update to exposure calculator (elevation included); discussion here initiated by Frazer on the missing antenna gains; Bryan has since sent out an update with plans to implement in the next go. OPT 1.05.04 notes below
      • 1. The OPT does not open any project until a user selects one.  Previously the OPT automatically opened the first project in the tree.  (For users whose first project is a large one, this has the happy side effect of allowing the OPT to open more quickly.)
         2. If User 1 has Project P open and User 2 attempts to view Project P, User 2 will be told that they are not allowed to even view the project until User 1 exits the OPT.  Previously User 2 could view, but not edit, Project P in this situation.  This is a stop-gap change that will probably be in place for about a month, at which time we hope to have a better, permanent, solution to the data-clobber problem.
         3. To get write (and read) access to Project P, User 2 no longer has to exit the OPT.  Once User 1 exits, User 2 may select Project P and begin to view and edit it.  Previously both Users 1 and 2 had to exit the OPT before write access could transfer from 1 to 2 for Project P.
        

Software status and updates (Butler)

  • Links:
  • Notes: Discussion: OBIT filler is undergoing significant testing; will this be made available through the archive; BB: after further testing is completed it can turn into an approved path. Anyone interested in testing this further please contact Eric locally.
  • CASA Update (new functionality available/issues/enhancements):JK: Release feature freeze has happened; bug fixes are ongoing. Note that progress will be tracked in JIRA; they've added an additional state="On Hold"; currently there are 600 open tickets of which only a fraction are underway so we may be seeing tickets updated in this way to clarify the development/work priorities. SM proposed another state called 'Deep Hold' that indicates things that are never planned to be worked on.
  • AIPS Update (new functionality available/issues/enhancements): Eric noted that the new public Lustre file system will take work on both the AIPS and user side to use; this is still underway. Working on sundry issues but noted that the 0900 Wednesday meeting has many details for those that are interested. Specific details can also be had through the AIPS Midnight Job - anyone who wants to be on this list please contact Eric (one note a day).

Observing recommendations update (Ops)

Commissioning priorities for the week (McMullin, Chandler)

  • 1) Software testing and updates (CBE, CM; work toward CBE averaging, work on missing/mis-labeled, robustness)
  • 2) Science observing as possible (Any array projects)
  • Interrupt status for additional 3-bit testing, 4-band tests needed.

Commissioning Reports

Correlator: Momjian: CBE Time Averaging Tests. https://bugs.nrao.edu/browse/ECSV-106

  • See the ticket for details. A discussion of the handling of flags was raised; this will be part of the final result. Currently this is advertised on the science pages and should be offered for C config; note that the 10s tests have been performed and analyzed and are consistent with expectations.

Spectral Line: Goss: Bandpass stability. https://bugs.nrao.edu/browse/ECSV-35

  • Preliminary results are shown in the 3 figures in the ticket. There are some antennas with issues (3-5); others are good. Test is ostensibly for viability of weak recombination/absorption line work (small fraction of the continuum); pre WIDAR would be able to get variations of 1 part in 1000. This study which compared 5 minute averages against the 8 hour integration shows a few parts in 10^4. Looking good; Laura will be out at the end of the week and they will assemble recommendations.

Sub GHz: Owen: 4-Band testing results

  • Six antennas had deployed dipoles; saw problems with 3 (much smaller BWs seen); a further test was taken and the BW approximation/shape have now been posted to the OSS.

In preparation

  • Polarization: Multi-band leakage/PA (Myers, Marvil): Not a priority but still needed; keep on list.
  • Correlator: Krauss/Hunter: 7 vs 8 node observing (follow-up; pending RSRO results)

Commissioning Test Planning

  • Correlator: CBE time averaging (>4 s)!
  • Archive: Update to MCAF - needs side-by-side comparison of the data model and results to deploy to production.
  • Calibration: Switched power tracking over several airmasses.
  • Calibration: Switched power on/off 1dB attenuators.
  • Polarization: Antenna cross polarization (Perley; Need script.); LSCX on 3C147; Delayed until Rick's return
  • Spectral Line: Dhawan: Follow-up on frequency offset phase jumps characterization.
  • Spectral Line: Ott: Doppler tracking sampling follow-up

RSRO/ECSO Reports

10B-126: Takahashi: Test run (TRSR0025) results. https://bugs.nrao.edu/browse/RSRO-31

  • 1 hour C band test looking to detect faint continuum sources. Strong HII region nearby so must mosaic a broader region to disentangle the emission from the sources of interest. Have begun making a test images of the region (3 fields at different distances, few arcminutes to 20 arcminutes).

10B-105: Fomalont: Test run (TRSR0028) results.

  • Ed is on travel but talked with him; struggling (as are others outside the building) to get and analyze the data. Saw some concerning issues with regard to the antenna motion for the mosaics (80s settle time for 1 deg separated points); still investigating. More later.
  • Coming

  • Notes:
    • Current Residents
      • Dave Green (POC: Sanjay Bhatnagar): area(s): algorithm tests for wideband imaging (currently away).
      • Volker Heesen (POC: Michael Rupen); area(s): polarimetry
      • Peter Hofner (POC: Juergen Ott); area(s): post-processing
      • Russ Taylor (POC: Rick Perley); area(s): polarimetry
      • Satoko Takahashi (POC: Juergen Ott); area(s): continuum, mosaicing; 10B-126 (AT380): A Millimeter Study of the Embedded Star-Cluster in the Orion Molecular Cloud.
        • Hauyu Liu (POC: Juergen Ott); area(s): continuum, mosaicing
      • Greg Hallinan (POC: Kumar Golap); area(s):; 10B-209 (AH1020): Broadband Periodic Dynamic Spectra of Ultracool Dwarf Pulsars
      • Adam Leroy (POC: TBD); area(s):; 10A-196 (AL746): Resolving the Starbursts in Nearby LIRGs and ULIRGs

Demo Science (Chandler/Condon): Project Codes: *TDEM*

Project Status Time Comments SB
Crab nebula Complete 10/10 L-band OSRO1/OSRO2 Complete
IRC+10216 Ready (HF) 16/33 Partial; revisit in DnC/C Partial
M87 Ready (HF) 0/6 Ka Not done
M51 Ready (MR) 8/8   Complete
Cygnus A Ready 3x6 1x10 possible; 12 OSRO hours obtained Withdrawn
M82 Complete 2/2 L band Complete
Herc A Ready 3/5 CX bands Partial

Commissioning Group updates

AOB

  • Claire noted that the PASEO report has come out; they've had a meeting discussing the various points; she has sent the report out to the group..

Misc Actions

  • [McMullin/Frail]: %REDWrite up disposition of RR items.
  • [McMullin]: Confirm restriction of RSRO scan length at 90 s
  • [Owen]: Outline issue with 90s scan length restriction.
  • [McMullin/Durand]: Clean up outdated MainSaver issues.
  • [McMullin]: Review/recover metrics for Priority 0 items.
  • [Perley/McMullin]: SPW/IF/BB lexicon and mapper
  • List of known problems (can be used by OSRO commissioning team (van Moorsel)
      • Med Priority [Rupen]: X, Ku band subbands are inverted (IF 1 is labelled as IF4, IF2 as 3, IF3 as 2, IF 4 as 1). Noted that this had been tested but there were enough changes to warrant a re-do.
      • Med Priority [TBD]: Phase offsets within scans (phase jump at scan boundaries). R&L phase offset on some scans - probably L302 LO problem, needs tracking down. MR: Seen on EA07, EA25 (both RR and LL)
      • Med Priority [TBD]: Added. Phase change when frequency changes (needs to be tested).
      • Med Priority [TBD]: Anti-aliasing does not behave as advertised.
      • Low Priority [TBD]: First scan often bad.
      • Low Priority [TBD]: Phase (and amp) wobbles; long history.
-- JosephMcMullin - 16 Sep 2010
Topic revision: r5 - 2010-09-22, jmcmulli
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding NRAO Public Wiki? Send feedback